OpenStudy (anonymous):

find the derivative of F(2)=(3/(x^2))

OpenStudy (anonymous):

Do you mean F(x)? \[ \frac{3}{x^2} = 3x^2 \] Try applying the power rule to that :)

OpenStudy (anonymous):

oops. I meant \[ \frac{3}{x^2} = 3x^{-2} \]

OpenStudy (anonymous):

Use the alternative form of the derivative to find the derivative of the function f(x)=(3/x^2) at x=2

OpenStudy (anonymous):

so it would be 3/4?

OpenStudy (anonymous):

It should be -3/4. But if they're asking you to use the alternative form, they're asking you to use the limit definition to evaluate it!

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

you're welcome :)